OpenStudy (anonymous):

Ok let's look at this example. (a) Approximate the function \[f(x)=\sqrt[3]{x}\] by a Taylor polynomial of degree 2 at a=8. \[f(x)=\sqrt[3]{x}\] \[f'(x)=\frac13 x^{-2/3}\] \[f''(x)=-\frac{2}{9} x^{-5/3}\] \[f'''(x)=\frac{10}{27} x^{-8/3}\] f(8)=2 \[f'(8)=\frac{1}{12}\] \[f''(8)=-\frac{1}{114}\] Thus: \[T_2(x)=f(8)+\frac{f'(8)}{1!}(x-8)+\frac{f''(8)}{2!}(x-8)^2\] so... \[\sqrt[3]{x}=T_2(x)=2+\frac{1}{12}(x-8)-\frac{1}{288}(x-8)^2\] This part makes sense, but what about this: (b) How accurate is this approximation when \[7\le x \le 9\]? Taylor inequality Theorem states \[f^{(n+

OpenStudy (anonymous):

I have to complete writing the second part. Give me a second

OpenStudy (anonymous):

The Taylor inequality Theorem States if \[\left|f^{(n+1)}(x)\right|\le M\] then \[\left|R_n(x) \right|\le \frac{M}{(n+1)!}{\left| x-1\right|}^{n+1}\]

OpenStudy (anonymous):

if n=2 and a=8 then \[\left|R_2(x) \right|\le \frac{M}{3!}{\left| x-8\right|}^{3}\]

OpenStudy (anonymous):

where \[f'''(x)\le M\] Because \[x\le 7\], we have \[x^{8/3} \le 7^{8/3}\] and so

OpenStudy (anonymous):

and so \[f'''(x)=\frac{10}{27}* \frac{1}{x^{8/3}} \le \frac{10}{27}*\frac{1}{7^{8/3}}<0.0021\]

OpenStudy (anonymous):

I don't understand part (b) of this question. Please please help

OpenStudy (anonymous):

Looks like you've already done it.

OpenStudy (anonymous):

It's an example from the book. I don't understand the example.... I can't do my homework if I don't even understand the example. I posted my hw question yesterday and no one could help me.

OpenStudy (anonymous):

It's example 1 in Chapter 11 section 11 of Stewart's Calculus.

OpenStudy (anonymous):

Oh, okay. Which part of it confuses you? For one thing, I think you typed the original equation incorrectly. \(\Large |R_n(x) |\le \frac{M}{(n+1)!}{\left| x-1\right|}^{n+1} \) should be \(\Large |R_n(x) |\le \frac{M}{(n+1)!}{\left| x-a\right|}^{n+1} \) right?

OpenStudy (anonymous):

Yes you are right. I guess I understand everything up to the 4th post but this part is still a mystery. \[f'''(x)=\frac{10}{27}* \frac{1}{x^{8/3}} \le \frac{10}{27}*\frac{1}{7^{8/3}}<0.0021\]

OpenStudy (anonymous):

hold on!!!!

OpenStudy (anonymous):

sry don't mean to scream but I think something just clicked

OpenStudy (anonymous):

so we replace M with f'''(x) in that inequality

OpenStudy (anonymous):

Right. Exactly. M comes from looking at the derivative one degree higher, so we look at f′′′(x)=(10/27)x^(−8/3) On the one side, at 7, we get some number, and as x increases, that M gets smaller and smaller. So we take the largest value of M because f''' will always be less than or equal to that M over the interval. Then we use M in the second equation.

OpenStudy (anonymous):

I didn't really state that very well, but we look at the n+1 derivative, and M is the greatest value of that derivative over the x interval.

OpenStudy (anonymous):

Let's step back for a moment and look at the bigger picture. I think I'm losing track of what the purpose of this inequality is.

OpenStudy (anonymous):

So we have this function f(x) and I'm trying to approximate that line by taking it's 2nd derivative. How am I doing so far?

OpenStudy (anonymous):

Well, you're trying to approximate the line by creating a polynomial. The polynomial uses the first derivative and the second derivative in this case. To make it more accurate, you can use the third derivative or even more.

OpenStudy (anonymous):

Understood. So when we say \[7 \le x \le9\] are we zooming in on a value between 7 and 9 (graphically speaking) 

OpenStudy (anonymous):

Right right. We're interested in the accuracy of our approximation between those two x values. We're not worried about how accurate it is when x is 10 or 10000

OpenStudy (anonymous):

ok. What is M? Just a value that is greater than next derivative up?

OpenStudy (anonymous):

Right. So if we look at \(\large f^{n+1}(x)\) over the interval and see that it's always less than or equal to some number, then we call that number M. We could use a number that's dramatically larger, but it's better if we use something that's the max value of the deriv or close to it.